OpenStudy (anonymous):

45y + x = 46 when x = 1

OpenStudy (anonymous):

y =1

OpenStudy (anonymous):

substitute 1 for x 45y + 1 = 46 subtract 1 45y + 1 -1 = 46 -1 45y = 45 divide by 45 45y/45 = 45/45 y = 1

OpenStudy (anonymous):

tbhanx